# The Role of Accuracy and Validation Effectiveness in Conversational Business Analytics

### v2

@@ 4.4 Comparison of Full Support with Partial Support and Delegation to the Data Engineer
The end user will choose FS only if it provides a higher expected value than the profit achievable through delegation
$E_{FS}(\alpha,\beta)>v$
and offers a better outcome than PS
@@
This condition (2) requires the validation to be sufficiently effective and robust (e.g., against the shortcomings of the end user conducting the validation based on the provided explanation) to compensate for inaccuracies in AI-generated information (($1-\alpha$)) and to deliver an overall expected value exceeding the guaranteed profit $v$. If the AI delegation condition for FS (2) is satisfied, it is prioritized over delegation to the data engineer. However, this preference also depends on the accuracy of AI-generated information being sufficiently high compared to the profit $v$. Specifically, the ”FS feasibility condition” must hold:
@@
This condition (3) ensures that the expected gain ($\alpha\cdot(+1)$) without validation exceeds the guaranteed profit $v$. If this condition is not satisfied, it is not justified to use FS, even with a perfect validation effectiveness, as the AI’s base accuracy would be insufficient to outperform human delegation. The accuracy threshold (3) for FS is lower than the threshold (1) for
@@
This indicates that the inclusion of a validation step in FS broadens the potential applicability of CBA. However, whether FS is advantageous compared to delegation to a data engineer depends on the level of validation effectiveness as defined in the AI delegation condition for FS (2).
@@
In condition (4), $\beta^{**}$ represents the threshold where validation effectiveness exceeds accuracy, making FS more advantageous than PS. It is important to note that all conditions described so far depend on the values of the potential gain ($+1$) and loss ($-1$). If the value of the loss outweighs that of the gain, the threshold for $\beta$ in the validation dominance condition shifts, allowing a less stringent validation standard for FS to outperform PS.
@@
Figure 6 illustrates the relationship between accuracy, validation effectiveness and the profit achievable through delegation to the data engineer (all values are measured in monetary units). The dotted area highlights the range of $\alpha$ and $\beta$ values where FS outperforms PS and the data engineer, as all three conditions regarding AI delegation (2), the feasibility of FS (3), and the dominance of validation (4) are satisfied. Outside this area, FS does not yield sufficient expected value to justify its use. The figure also identifies the threshold $\alpha^{*}_{PS}$ which is determined by the AI delegation condition for PS (1).
@@
In region $A$, the FS feasibility condition (3) is not met, as the accuracy of information generation is less than or equal to the profit achievable through the data engineer ($\alpha\leq v$). In this region, neither PS nor FS offers sufficient expected value. PS fails due to low accuracy, and FS is unable to compensate for this limitation, even when validation effectiveness would be perfect. Delegation to the data engineer remains the only viable option for retrieving the KPI. In this scenario, CBA is not utilized because it fails to match the effectiveness of the data engineer, rendering CBA unsuccessful. Furthermore, if the data engineer provides the KPI with significant delay, rendering it irrelevant for the end user due to its diminished value for timely decision-making ($v$ approaching zero), this not only signifies the failure of CBA but also highlights a broader limitation of business analytics in fulfilling its fundamental objective: delivering timely and decision-relevant information.
@@
Moreover, the requirements for accuracy in PS and FS, as well as for validation effectiveness in FS, increase as $v$ rises. A reduction in urgency causes a rightward shift of $v$, $\alpha^{*}_{PS}$, and the delegation threshold condition for FS (2). In other words, the less urgent the KPI (making it less problematic to wait for the data engineer) the greater the performance demands on CBA to generate the correct SQL query and support the reliable assessment of its correctness.
@@
In Region $B$, the FS feasibility condition (3) is met ($\alpha>v$), but the AI delegation condition for PS (1) is not satisfied ($\alpha\leq\alpha^{*}_{PS}$). In this region, PS is less effective than delegating to the data engineer. However, FS becomes a viable option if validation effectiveness meets the AI delegation condition for FS (2). Under this condition, validation introduces a ”boosting effect” by identifying and mitigating errors in SQL queries, ensuring that flawed KPIs do not influence the decision-making process. This boosting effect enables FS to enhance the overall reliability of information, compensating for the limitations of lower accuracy. For example, even if accuracy falls below the threshold required for PS (1) (e.g., $\alpha_{1}$), sufficiently high validation effectiveness in FS (e.g., $\beta_{1}$) can still yield an expected value greater than $v$. While the generation of the SQL query may fail, resulting in incorrect KPI, sufficiently effective validation can still identify these errors, ensuring that incorrect KPIs do not adversely affect the decision process.
@@
In region $C$, where the AI delegation condition for PS (1) is met ($\alpha>\alpha^{*}_{PS}$), PS becomes a viable alternative to delegation to the data engineer, as it provides sufficient benefit on its own. If the AI delegation condition for FS (2) and the validation dominance condition (4) are not fulfilled (e.g., at $\alpha_{2}$ and $\beta_{2}$), FS has a ”devastating effect” as the validation process undermines the reliability of the SQL queries generated. Not only are erroneous SQL queries less effectively identified, but insufficient validation also leads to the unwarranted rejection of correct SQL queries. Consequently, FS is not only less effective than PS but also inferior to delegation to the data engineer. Therefore, using AI-based assistance for KPI generation with AI-supported validation is not recommended. This underscores the inherent risks of relying on user-based validation without robust mechanisms to support in correctly evaluating AI-generated SQL queries. In cases where expertise is low or explanations are unclear, FS may inadvertently introduce errors or misjudgments.
@@
FS becomes favorable when, in region $C$, both conditions, (2) and (4), are fulfilled (e.g., at $\beta_{2}$ and $\alpha_{3}$). In such cases, the validation process enhances the reliability of the generated KPI and maximizes the benefits of FS. As accuracy increases, the requirements for validation effectiveness must also rise to ensure that FS remains beneficial. A particular challenge arises when accuracy is very high; in such cases, user-based validation must become even more effective to provide added value over PS. This places significant demands on validation mechanisms, especially in terms of users’ ability to evaluate and assess complex SQL queries based on the given explanations.
These findings highlight the critical interplay between accuracy and validation effectiveness in determining the optimal use of PS or FS in CBA. High accuracy favors PS, while lower accuracy requires robust validation effectiveness for FS to deliver reliable results. Insufficient validation, however, compromises performance, making delegation to a data engineer the preferable alternative. This analysis emphasizes the importance of aligning validation performance with accuracy to enhance the effectiveness of CBA. Although FS can improve decision-making, its success depends on validation mechanisms that appropriately scale with accuracy. When this balance cannot be achieved, PS or data engineer may offer more dependable outcomes.
@@ 5 Conclusion
This study analyzes CBA, an emerging approach with considerable potential to address the competency gaps inherent in traditional self-service analytics. Through the use of LLMs, CBA facilitates natural language interactions, enabling end users to independently perform tasks such as data retrieval, preparation, and insight generation. In this context, CBA represents a novel approach to task delegation in business analytics, shifting the process of information generation from human experts to AI, thereby making advanced analytics more timely and accessible to a broader range of organizational roles. A focus of this study is Text-to-SQL, a well-established semantic parsing technology that has seen transformative advancements through the application of LLMs, enhancing its ability to translate natural language requests into structured SQL statements.
@@
The theoretical models developed in this study examine the conditions under delegation to AI, through PS and FS, outperforms the delegation of tasks to human experts. PS is effective when the accuracy ($\alpha$) exceeds a specified threshold, as outlined in condition (1). FS, however, is only beneficial when three conditions are met: the AI delegation condition (2), FS feasibility condition (3), and validation dominance condition (4).
@@
The study highlights the critical importance of validation effectiveness ($\beta$) in the context of FS. FS can exert a "boosting effect" on information generation, particularly when validation is robust. Even with low accuracy ( $\alpha$), a positive benefit can be derived from the delegation process by preventing incorrect information from being incorporated into operational decision-making. However, the reliance on user-based validation introduces significant uncertainties, especially when users lack the necessary expertise or when SQL queries and data models are characterized by high complexity. These challenges can have a ”devastating effect” on the quality of information generation, potentially making FS less effective than PS and, in the worst-case scenario, even less effective than delegation to a data engineer. This underscores the need for robust validation mechanisms that support in correctly evaluating AI-generated information. Without such mechanisms, the advantages of CBA – particularly when the accuracy ($\alpha$ ) of translating natural language requests into correct information is high – could be significantly diminished.
@@
A promising avenue for future research is the development of hybrid approaches that combine existing techniques (e.g., decomposition, visualization and dialogue techniques) to support code comprehension with automated validation techniques. Future studies could explore the implementation of automated methods for user-independent assessment of the reliability of generated KPIs. Potential approaches might include statistical methods, such as confidence intervals to evaluate the correctness of KPIs, and cross-validation with previously validated reports and KPIs. Integrating such methods could mitigate the risk of misclassification due to a lack of expertise, thereby enhancing the robustness of CBA.
@@
The theoretical models presented in this study are based on several simplifying assumptions. First, the gain from making the correct decision is fixed at $+1$, and the loss from making the wrong decision is set to $-1$. While these values determine the derived conditions, they may not fully capture the variability observed in real-world scenarios. Furthermore, the analysis adopts a static perspective, assuming a single execution cycle for KPI creation and validation, thereby overlooking the iterative nature of CBA. The models also disregard the costs associated with using LLMs, data storage, and processing, potentially leading to an overestimation of the economic feasibility of CBA. Additionally, they assume perfect data quality, which may not align with practical conditions. Extending these models to incorporate dynamic, cost-sensitive, multi-stage decision-making processes and data quality considerations could yield deeper and more realistic insights.
In summary, this study highlights the transformative potential of CBA in overcoming the limitations of traditional self-service analytics and advancing the capabilities of business analytics. However, to fully realize this potential, it is crucial to address inherent challenges related to accuracy and validation, while exploring the added value of automated validation techniques. Integrating CBA into business analytics can improve data-driven decision-making and enhance accessibility to advanced analytics. With continued innovation, CBA holds the promise of becoming a trusted and indispensable approach for informed decision-making across diverse business contexts.

### v3

@@ 4.4 Comparison of Full Support with Partial Support and Delegation to the Data Engineer
The end user will choose FS only if it provides a higher expected value than the profit achievable through delegation,
$E_{FS}(\alpha,\beta)>v,$
and offers a better outcome than PS,
@@
Condition (2) requires the validation to be sufficiently effective and robust (e.g., against the shortcomings of the end user conducting the validation based on the provided explanation) to compensate for inaccuracies in the AI-generated information (($1-\alpha$)) and to deliver an overall expected value exceeding the guaranteed profit $v$. If the AI delegation condition for FS (2) is satisfied, it is prioritized over delegation to the data engineer. However, this preference also depends on the accuracy of AI-generated information being sufficiently high compared to the profit $v$. Specifically, the ”FS feasibility condition” must hold:
@@
Condition (3) ensures that the expected gain ($\alpha\cdot(+1)$) without validation exceeds the guaranteed profit, $v$. If this condition is not satisfied, it is not justified to use FS, even with perfect validation effectiveness, as the AI’s base accuracy would be insufficient to outperform human delegation. The accuracy threshold (3) for FS is lower than the threshold (1) for
@@
This indicates that the inclusion of a validation step in FS broadens the potential applicability of CBA. However, whether FS is advantageous compared with delegation to a data engineer depends on the level of validation effectiveness, as defined in the AI delegation condition for FS (2).
@@
All $\beta$ values that satisfy (4) make FS more advantageous than PS. It is important to note that all the conditions described so far depend on the values of the potential gain ($+1$) and loss ($-1$). If the value of the loss outweighs that of the gain, the threshold for $\beta$ in the validation dominance condition shifts, allowing a less stringent validation standard for FS to outperform PS.
@@
Fig. 6 illustrates the relationship between accuracy, validation effectiveness and the profit achievable through delegation to the data engineer (all values are measured in monetary units). The dotted area highlights the range of $\alpha$ and $\beta$ values where FS outperforms PS and the data engineer, as all three conditions regarding AI delegation (2), feasibility of FS (3), and dominance of validation (4) are satisfied. Outside this area, FS does not yield sufficient expected value to justify its use. The figure also identifies the threshold $\alpha^{*}_{PS}$ which is determined by the AI delegation condition for PS (1).
@@
In region $A$, the FS feasibility condition (3) is not met, because the accuracy of information generation is less than or equal to the profit achievable through the data engineer ($\alpha\leq v$). In this region, neither PS nor FS offer sufficient expected values. PS fails owing to its low accuracy, and FS is unable to compensate for this limitation, even when validation effectiveness would be perfect. Delegation to the data engineer remains the only viable option for retrieving the KPI. In this region, CBA is not utilized because it fails to match the effectiveness of the data engineer, rendering it unsuccessful. Furthermore, if the data engineer provides the KPI with significant delay, rendering it irrelevant for the end user due to its diminished value for timely decision-making ($v$ approaching zero), this not only signifies the failure of CBA but also highlights a broader limitation of business analytics in fulfilling its fundamental objective of delivering timely and decision-relevant information.
@@
Moreover, the requirements for accuracy in PS and FS, as well as for validation effectiveness in FS, increase as $v$ rises. A reduction in urgency causes a rightward shift of $v$, $\alpha^{*}_{PS}$, and the delegation threshold condition for FS (2). In other words, the less urgent the KPI (making it less problematic to wait for the data engineer), the greater the performance demands on CBA to generate the correct SQL query and support the reliable assessment of its correctness.
@@
In Region $B$, FS feasibility condition (3) is met ($\alpha>v$), but the AI delegation condition for PS (1) is not satisfied ($\alpha\leq\alpha^{*}_{PS}$). In this region, PS is less effective than delegating to the data engineer. However, FS becomes a viable option if the validation effectiveness satisfies the AI delegation condition for FS (2). Under this condition, validation introduces a ”boosting effect” by identifying and mitigating errors in SQL queries, ensuring that flawed KPIs do not influence the decision-making process. This boosting effect enables FS to enhance the overall reliability of the information, compensating for the limitations of lower accuracy. For example, even if the accuracy falls below the threshold required for PS (1) (e.g., $\alpha_{1}$), a sufficiently high validation effectiveness in FS (e.g., $\beta_{1}$) can still yield an expected value greater than $v$. While the generation of the SQL query may fail, resulting in incorrect KPI, sufficiently effective validation can still identify these errors, ensuring that incorrect KPIs do not adversely affect the decision process.
@@
In region $C$, where the AI delegation condition for PS (1) is met ($\alpha>\alpha^{*}_{PS}$), PS becomes a viable alternative to delegation to the data engineer as it provides sufficient benefit on its own. If the AI delegation condition for FS (2) and the validation dominance condition (4) are not fulfilled (e.g., at $\alpha_{2}$ and $\beta_{2}$), FS has a ”devastating effect” because the validation process undermines the reliability of the SQL queries generated. Not only are erroneous SQL queries less effectively identified, but insufficient validation also leads to the unwarranted rejection of correct SQL queries. Consequently, FS is not only less effective than PS but also inferior to delegation to the data engineer. Therefore, AI-based assistance for KPI generation with AI-supported validation is not recommended. This underscores the inherent risk of relying on user-based validation without robust mechanisms to support the correctly evaluation of AI-generated SQL queries. In cases where expertise is low or explanations are unclear, FS may inadvertently introduce errors or misjudgments.
@@
FS becomes favorable when, in region $C$, conditions (2) and (4) are fulfilled (e.g., at $\beta_{2}$ and $\alpha_{3}$). In such cases, the validation process enhances the reliability of the generated KPI and maximizes the benefits of FS. As the accuracy increases, the requirements for validation effectiveness must also increase to ensure that FS remains beneficial. A particular challenge arises when the accuracy is very high; in such cases, user-based validation must become even more effective to provide added value over PS. This places significant demands on validation mechanisms, particularly in terms of users’ ability to evaluate and assess complex SQL queries based on the given explanations.
@@ 5 Conclusion
This study analyzes CBA, an emerging approach with considerable potential to address the skill gaps inherent in traditional self-service analytics. Through the use of LLMs, CBA facilitates natural language interactions, enabling end users to independently perform tasks such as data retrieval, preparation, and insight generation. In this context, CBA represents a novel approach to task delegation in business analytics, shifting the process of information generation from human experts to AI, thereby making advanced analytics more timely and accessible to a broader range of organizational roles. This study focuses on Text-to-SQL, a semantic parsing technology that has seen transformative advancements through the application of LLMs, enhancing its ability to translate natural language requests into structured SQL statements.
@@
The theoretical models developed in this study examine the conditions under delegation to AI, through PS and FS, outperforms the delegation of tasks to human experts. PS is effective when the accuracy ($\alpha$) exceeds a specified threshold, as outlined in condition (1). FS, however, is only beneficial when three conditions are met: AI delegation condition (2), FS feasibility condition (3), and validation dominance condition (4).
@@
This study highlights the importance of validation effectiveness ($\beta$) in the context of FS. FS can exert a ”boosting effect” on information generation, particularly when the validation is robust. Even with low accuracy ( $\alpha$), a positive benefit can be derived from the delegation process by preventing incorrect information from being incorporated into operational decision-making. However, the reliance on user-based validation introduces significant uncertainties, especially when users lack the necessary expertise or when SQL queries and data models are highly complex. These challenges can have a ”devastating effect” on the quality of information generation, potentially making FS less effective than PS and, in the worst-case scenario, even less effective than delegation to a data engineer. This underscores the need for robust validation mechanisms that support in correctly evaluating AI-generated information. Without such mechanisms, the advantages of CBA – particularly when the accuracy ($\alpha$ ) of translating natural language requests into correct information is high – could be significantly diminished.
@@
A promising avenue for future research is the development of hybrid approaches that combine existing techniques (e.g., decomposition, visualization and dialogue) to support code comprehension with automated validation techniques. Future studies could explore the implementation of automated methods for user-independent assessment of the reliability of the generated KPIs. Potential approaches might include statistical methods, such as confidence intervals, to evaluate the correctness of KPIs, and cross-validation with previously validated reports and KPIs. Integrating such methods could mitigate the risk of misclassification owing to a lack of expertise, thereby enhancing the robustness of CBA.
@@
The theoretical models presented in this study are based on several simplifying assumptions. First, the gain from making the correct decision is fixed at $+1$ and the loss from making the wrong decision is set to $-1$. Although these values determine the derived conditions, they may not fully capture the variability observed in real-world scenarios. Furthermore, the analysis adopts a static perspective, assuming a single execution cycle for KPI creation and validation, thereby overlooking the iterative nature of CBA. The models also disregard the costs associated with using LLMs, data storage, and processing, potentially leading to overestimation of the economic feasibility of CBA. Additionally, they assume perfect data quality, which may not align with practical conditions. Extending these models to incorporate dynamic, cost-sensitive approaches to information generation and account for data quality considerations could provide deeper and more realistic understanding.